Please show all work.

FINA Inc. considers a project with the following information:

Initial Outlay: 1,500

After-tax cash flows:

Year 1: -$100

Year 2: $1000

Year 3: $700

FINA’s assets are $500 million, financed through bank loans, bonds, preferred stocks, and common stocks. The amounts are as follows: Bank loans: $ 100 million borrowed at 10% Bonds: $180 million, paying 9% coupon with quarterly payments, and maturity of 5 years. FINA sold its $1,000 par-value bonds for $1,070 and had to incur $20 flotation cost per bond. Preferred Stocks: $20 million, paying $15 dividends per share. FINA sold its preferred shares for $210 and had to incur a $10/share flotation cost. Common Stocks: $200 million, the beta of FINA stocks is 1.5, the 90 day Treasury yield is 5%, and the return on the market portfolio is 15 %. FINA is subject to a 20% tax rate. Assuming the company uses WACC to compute the present value of the future cash flows, please find the following:

10) FINA is expected to pay a $4 per share common stock dividend at the end of this year. The dividends are expected to grow at 6% per year forever. How much should be the value of FINA’s shares?
Answer #1

10). Cost of equity (ke) = risk-free rate + beta*(market return - risk-free rate)

= 5% + 1.5*(15%-5%) = 20%

Intrinsic price per share = expected dividend/(ke - perpetual growth rate)

= 4/(20%-6%) = 28.57 per share